Opened 10 years ago

Closed 9 years ago

Last modified 2 years ago

#268 closed flight processing (fixed)

GB07/07, flight day 277, Inverclyde

Reported by: mark1 Owned by: knpa
Priority: alpha 4 high Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by knpa)

Data location: ~arsf/arsf_data/2009/flight_data/uk/GB07_07-2009_277_Inverclyde

Data arrived from ARSF via SATA disk L on 7th October 2009.

Scientific objective: Effect on spectral reflectance by heather type and biophysical properties

Priority: alpha-4-High

PI: Alasdair MacArthur

Sensors:

  • Eagle (07/01/2010)
  • Hawk (07/01/2010)
  • Leica LIDAR (27/04/2010)

Change History (34)

comment:1 Changed 10 years ago by mark1

  • Description modified (diff)

Still awaiting flight logsheet. When it arrives check eagle flight lines are correct - they are named 274 not 277 (date stamps match those of hawk though).

DEM creation gave errors in Grass but output a DEM. Check its good when processing.

comment:2 Changed 10 years ago by mark1

  • Priority changed from immediate to alpha 4 high

comment:3 Changed 10 years ago by mark1

logsheet arrived and unpacked

comment:4 Changed 10 years ago by mark1

Have renamed eagle lines from VNIR274-09 to VNIR277-09, and from VNIRdatacube to VNIR274-09-1

comment:5 Changed 10 years ago by mark1

Have renamed Hawk lines from Hawk277* to SWIR277*

comment:6 Changed 10 years ago by mggr

Re: #281 (Hawk frame rates are double what they should be), the following lines are affected and the .hdr file will need to be edited to halve the fps number:

uk/GB07_07-2009_277_Inverclyde/hawk/SWIR277-09-1.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB07_07-2009_277_Inverclyde/hawk/SWIR277-09-2.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB07_07-2009_277_Inverclyde/hawk/SWIR277-09rec-1.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB07_07-2009_277_Inverclyde/hawk/SWIR277-09rec-2.hdr: predicted length half recorded length, frame rate doubling problem?

comment:7 Changed 10 years ago by harg

  • Owner set to harg

comment:8 Changed 10 years ago by chrfi

hawk flight line 3 no .hdr file

comment:9 Changed 10 years ago by chrfi

  • Owner changed from harg to chrfi
  • Status changed from new to assigned

comment:10 Changed 10 years ago by chrfi

hawk line 3 lacking dark frames, used dark frames from line 2

comment:11 Changed 10 years ago by chrfi

last 111 lines of eagle 3 were removed to allow the line to process

comment:12 Changed 10 years ago by chrfi

eagle

linesctgpt
1-0.02-0.05
20.040.00
30.000.00
4-0.0650.04
5-0.030.02

hawk

linesctgpt
11.600.00
20.820.06
377.70.10
418.650.00
50.100.00

comment:13 Changed 10 years ago by benj

  • Other processors set to benj

Doing delivery check

comment:14 Changed 10 years ago by benj

Updated copyright notice to 2010, a few minor readme changes, removed superfluous readme template file. Two of the Hawk files appear to have the wrong filenames (should be fixed before delivery).

Data values in Hawk line 3 are very low - underflows in a lot of bands, which means probably the values in the rest of the bands are wrong. This should be mentioned in the readme. Also the cause should be checked, may just have been that the sensor was having problems on line 3 (hence why the end had to be cut off).

Otherwise ready for delivery.

comment:15 Changed 10 years ago by benj

Low data values in Hawk line 3 appear to be a result of the dark frame substitution - substituting dark frames on another line on the same flight causes the same problem. Suggests that the data values on all of the Hawk lines are suspect and shouldn't be relied on to be correct.

comment:16 Changed 10 years ago by chrfi

hyperspectral data delivered 07/01/2010

comment:17 Changed 10 years ago by chrfi

files in workspace need to be rsynced back to arsf_data as soon as space is available

comment:18 Changed 10 years ago by benj

There's now space on the disk, you can rsync this back.

comment:19 Changed 9 years ago by emca

LiDAR delivery folder created by chrfi at
~airborne/workspace/GB07_07-2009_277_Inverclyde/delivery/20100426/GB07_07

Changed PI on logsheet & readme to Alasdair MacArthur.
Removed .xml file from DEM directory.

Accuracy to vectors needs to be added to read_me. Once that is added the data is ready to deliver.

comment:20 Changed 9 years ago by chrfi

lidar roll offsets

115948 0.003903
115218 0.003903
121017 -0.003663
123771 -0.00343
122647 -0.004

comment:21 Changed 9 years ago by chrfi

  • Resolution set to fixed
  • Status changed from assigned to closed

comment:22 Changed 9 years ago by chrfi

  • Other processors benj deleted
  • Resolution fixed deleted
  • Status changed from closed to reopened

my mistake, afaik it needs archiving

comment:23 Changed 9 years ago by chrfi

  • Owner changed from chrfi to knpa
  • Status changed from reopened to assigned

comment:24 Changed 9 years ago by chrfi

  • Component changed from Processing: general to Archiving
  • Owner changed from knpa to chrfi

comment:25 Changed 9 years ago by chrfi

waiting for more space before rsyncing back

comment:26 Changed 9 years ago by knpa

  • Description modified (diff)
  • Owner changed from chrfi to knpa

Beginning archiving.

comment:27 Changed 9 years ago by knpa

No webcam data for this project.

comment:28 Changed 9 years ago by knpa

No az run logs present.

comment:29 Changed 9 years ago by knpa

Hawk raws have odd names.
1-3 OK then assume SWIR277-09rec-1 is 4 and SWIR277-09rec-2 is 5.

comment:30 Changed 9 years ago by knpa

Archiving preparation completed, awaiting NEODC transfer.

comment:31 Changed 9 years ago by knpa

Downloaded and backed up by NEODC.

Moved to archive disk.

comment:32 Changed 9 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace version.

Closing ticket.

comment:33 Changed 5 years ago by dap

Delivery checking reprocessed data for delivery to PI:

  • Copied data quality report to doc/
  • Copied xml/css stylesheets to project_information and line_information
  • Generated ASCII view vectors files for Eagle and Hawk sensors and converted line endings to CRLF.

Delivering dataset to PI via FTP.

comment:34 Changed 2 years ago by asm

Archiving

Archived and it is available from CEDA. Re-processed hyperspectral data at http://browse.ceda.ac.uk/browse/neodc/arsf/2009/GB07_07/GB07_07-2009_277_Inverclyde/hyperspectral-reprocessed

Plymouth 12th May 2017

Note: See TracTickets for help on using tickets.